OpenStudy (anonymous):

Perform the operation; (3.9 -(3)^2) divide (8.3 -8)x4+(-1)

OpenStudy (anonymous):

you need to follow order of operation, PEMDAS Please Excuse My Dear Aunt Sally: v Highest Priority P arentheses E xponents M ultiplication and D ivision, A ddition and S ubtraction ^ Lowest Priority

OpenStudy (anonymous):

i know can you solve it

OpenStudy (anonymous):

part of the problem is: -(3)^2 you want to try ?

OpenStudy (anonymous):

what does it simplify to

OpenStudy (anonymous):

-(3)^2

OpenStudy (anonymous):

can you just do it in steps

OpenStudy (anonymous):

yes: -(3)^2 the exponent is linked to the parentheses, so it is linked to the number in parentheses the number in parentheses is "3" only, so we get after doing the exponent: with (3)^2 = 9: - 9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I've started with the left parentheses and looked for any parentheses, and the (3)^2 was the last remaining one on this "level" (3.9 -(3)^2) divide (8.3 -8)x4+(-1) became (3.9 - 9 ) divide (8.3 -8)x4+(-1)

OpenStudy (anonymous):

now there is no longer a parentheses inside of another parentheses, so the next step is to follow order of operations in each of the remaining par.

OpenStudy (anonymous):

(3.9 - 9 ) divide (8.3 -8)x4+(-1) we got three of them: 1. (3.9 - 9) 2. (8.3 - 8) 3. (-1)

OpenStudy (anonymous):

3.9 - 9 so that is like 3.0 - 9 = -6+0.9 = -5.1

OpenStudy (anonymous):

i first subtract 9 and then add back .9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok, next one is (8.3 - 8) is easier, that will be 0.3 right? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

finally 3. (-1) is just -1 we get

OpenStudy (anonymous):

(-5.1 ) divide (0.3)x4+-1

OpenStudy (anonymous):

also, 3+-2 is the same as 3-2 (-5.1 ) divide (0.3)x4-1

OpenStudy (anonymous):

I believe this is the denominator in the original problem: (8.3 -8)x4+(-1) it makes a difference for the +1 term, if the addition is part of the denominator, then there is like an invisible parentheses that is around the entire denominator

OpenStudy (anonymous):

The problem was like this? \[\frac{ (3.9 -(3)^2) }{ (8.3 -8)x4+(-1) }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so if there is an addition in denominator thats the same like parentheses around the denominator this is equivalent: \[(3.9 -(3)^2) \div [(8.3 -8)x4+(-1)]\]

OpenStudy (anonymous):

that means we must first simplify the denominator completely before dividing

OpenStudy (anonymous):

last result was: (-5.1 ) divide (0.3)x4-1 with fraction:\[\frac{ (-5.1 ) }{ (0.3)x4-1 }\]

OpenStudy (anonymous):

next task: simplify [(0.3)x4-1]

OpenStudy (anonymous):

multiplication comes before subtraction, 0.3 x 4 = 1.2

OpenStudy (anonymous):

result [1.2-1]

OpenStudy (anonymous):

further simlify: 1.2 - 1 = 0.2

OpenStudy (anonymous):

this gives \[\frac{ (-5.1 ) }{ 0.2 }\]

OpenStudy (anonymous):

it has to be -69

OpenStudy (anonymous):

we have made an error.... this one gets to -25,5

OpenStudy (anonymous):

(3.9 -(3)^2) divide (8.3 -8)x4+(-1)

OpenStudy (anonymous):

so what is the right answer

OpenStudy (anonymous):

(3.9 -9) divide (8.3 -8)x4+(-1) (-5.1) divide (0.3)x4+(-1) (-5.1) divide 1.2+(-1) (-5.1) divide 0.2 =-25.5

OpenStudy (anonymous):

ok thanks can u help me with one more

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7/12, 1 5/8 and 4 5/6 Find the average

OpenStudy (anonymous):

average is adding all of them then dividing through count of values

OpenStudy (anonymous):

to add them they need a common denominator

OpenStudy (anonymous):

LCF: 12, 8, 6

OpenStudy (anonymous):

7/12, 1 5/8 and 4 5/6

OpenStudy (anonymous):

we could make denominator a 24

OpenStudy (anonymous):

7/12, 1 5/8 and 4 5/6 will become 14/24, 1 15/24 and 4 20/24

OpenStudy (anonymous):

i extended first fractoin with 2 (2x12=24), second fraction with 3 (8x3=24) and third fraction with 4 (6x4=24) then they all have the same denominator

OpenStudy (anonymous):

ok yea i got it

OpenStudy (anonymous):

1 15/24 is the same as 24/24 + 15/24 39/24

OpenStudy (anonymous):

14/24 + 39/24+ "4" 20/24 4 is the same as 4*24/24 or 96/24

OpenStudy (anonymous):

14/24 + 39/24+ 96/24+20/24

OpenStudy (anonymous):

=14/24 + 39/24+ 106/24+10/24 =14/24 + 39/24+ 116/24

OpenStudy (anonymous):

14+16 = 20+10 = 30 so 14/24 + 116/24 = 130/24

OpenStudy (anonymous):

total value of all values: 39/24 + 130/24

OpenStudy (anonymous):

39/24 + 130/24 = 9/24 + 160/24 = 169/24

OpenStudy (anonymous):

that is the total value, to get average divide by count of values whats the count of values in our example?

OpenStudy (anonymous):

i was supposed to get 2 25/72

OpenStudy (anonymous):

72 is three times 24

OpenStudy (anonymous):

ill try again 7/12, 1 5/8 and 4 5/6

OpenStudy (anonymous):

14/24, 13/8 and 29/6

OpenStudy (anonymous):

14/24, 39/24 and 116/24

OpenStudy (anonymous):

169/24

OpenStudy (anonymous):

then a third of that

OpenStudy (anonymous):

either we divide the numerator by three or we multiply the denominator with three, has the same effect

OpenStudy (anonymous):

since we can not evenly divide 169 by 3 (result: 56,333333) maybe its better to multiply the denominator then

OpenStudy (anonymous):

169/72

OpenStudy (anonymous):

is a third of the total value, three because three values

OpenStudy (anonymous):

169/72 can be simplified since 169 is more than 72

OpenStudy (anonymous):

it is two times 72 and then some 2x72 = 144

OpenStudy (anonymous):

so we have 2 REST/72 the rest is 169-144(the whole 2) = 25

OpenStudy (anonymous):

so 2 25/72

OpenStudy (anonymous):

they got this result by multiplying the denominator when we had to divide the fraction by three :) divide value of fraction by three is the same as multiplying its denominator by three

OpenStudy (anonymous):

what about this one: 10 - 4 5/8

OpenStudy (anonymous):

the first one is a number and the second one number + fraction

OpenStudy (anonymous):

4 5/8 means it has 4 times a fully 8/8 fraction then also a 5/8 fraction: 4 5/8 is 8/8 + 8/8 + 8/8 + 8/8 + 5/8

OpenStudy (anonymous):

so, it is 32/8 + 5/8 = 37/8

OpenStudy (anonymous):

10 is the same as 10/1 and that can be extended by 8 up and down: \[10 = \frac{ 10 }{ 1 }\]\[\frac{ 10 \times 8 }{ 1 \times 8 } = \frac{ 80 }{ ? }\]

OpenStudy (anonymous):

\[= \frac{ 80 }{ 8 }\]

OpenStudy (anonymous):

if they have the same denominator, we can add/subtract fractions

OpenStudy (anonymous):

\[\frac{ 80 }{ 8 } - \frac{ 37 }{ 8 }\]

OpenStudy (anonymous):

this is equivalent to 10 - 4 5/8

OpenStudy (anonymous):

\[\frac{ 80 }{ 8 } - \frac{ 37 }{ 8 }\] is equivalent to the problem 10- 4 5/8

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no problem